# Voltage-tuned anomalous-metal to metal transition in hybrid Josephson junction arrays

### v1

@@ Abstract
The measured hybrid Josephson junction arrays were fabricated on an InAs two-dimensional electron gas (2DEG) heterostructure proximitized by epitaxial Al, similar to that described in Refs. [37, 38]. The Al layer was lithographically patterned into a $255\times 31$ array of $1\,\mu$m square islands with 400 nm spacing. After covering the structure with a thin HfOx dielectric, a 100 nm wide frame gate was metallized between the Al islands to tune their coupling; see Fig. 1(a). A second HfOx layer and a global Ti/Au top gate were added to control the uncovered regions around the islands. A schematic cross-section of the complete device stack is shown in Fig. 1(b). An optical image of the final device, overlaid with the measurement setup is shown in Fig. 1(c). Two additional devices, with different island sizes and spacings, showed similar results. In the main text, we focus on data from a representative device. Supporting data from the other two devices are presented in the Supplementary Material [39]. Measurements were performed using standard lock-in techniques in a dilution refrigerator with a three-axis vector magnet and base temperature of 15 mK. Further details on the growth, sample fabrication, and measurements are provided in the Supplemental Material [39].
@@
Figure 3: (a) Resistivity, $\rho$, as a function of temperature, $T$, measured for different frame-gate voltage, $V_{\mathrm{FG}}$, values ranging from $-1.060$ to $-1.225$ V, at fixed $V_{\mathrm{TG}}=0$. The dashed black lines are linear fits of $\ln(\rho)$ versus $\ln(T)$ up to $70$ mK. (b) Low-$T$ resistivity slope, $d\ln(\rho)/d\ln(T)$, taken from the fits in (a), plotted against the average resistivity, $\bar{\rho}$, in the same $T$ range. Error bars are the standard uncertainties of the fitting.
@@
We note that $\rho$ at $V_{\mathrm{FG}}=-1.20$ V remains roughly $h/4e^{2}$ throughout the measured $T$ range [Fig. 3(a)]. For less negative $V_{\mathrm{FG}}$, $\rho$ drops below its normal-state value—in some cases by several orders of magnitude—before saturating at low $T$. For slightly more negative $V_{\mathrm{FG}}$, the first couple of traces still saturate, but at values higher than their high-$T$ resistivity. This behavior resembles a smooth crossover from a failed superconductor to a failed insulator within the anomalous metallic phase and may be attributed to quantum fluctuations [41, 34, 42].
@@
Figure 4: (a) Average resistivity, $\bar{\rho}$, measured as a function of top-gate, $V_{\mathrm{TG}}$, and frame-gate, $V_{\mathrm{FG}}$, voltages, showing an enlargment of the region where the superconducting, insulating, and metallic states meet. The plot was constructed by averaging a set of $\rho$ maps over the temperature range from $T=15$ to 55 mK, taken every $2.5$ mK. (b) Low-$T$ $\rho$ slope, $d\ln(\rho)/d\ln(T)$, obtained by linear fitting of $\ln(\rho)$ versus $\ln(T)$ for the same data as in (a)
@@
All data discussed so far were taken at magnetic field carefully tuned to zero (for more details, see Supplemental Material [39]). We now turn to investigate the effect of an out-of-plane magnetic field, $B$, which introduces a superconducting phase gradient between neighboring islands [43]. To this end, we map the resistivity as a function of $T$ and magnetic frustration, $f=\Phi/\Phi_{0}$, where $\Phi=BA$ is the magnetic flux through a unit cell of area $A$ and $\Phi_{0}=h/2e$ is the superconducting flux quantum. We find that the critical temperature, $T_{\mathrm{C}}$, is maximal at $f=0$ and decays quickly as the array is frustrated; see Fig. 5(a). Additional peaks in $T_{\mathrm{C}}$ appear at several other frustration values, most prominently at $f=1/2$, corresponding to commensurate vortex configurations.
@@
In both cases, the emergence of a small but finite positive slope at nonzero $f$ indicates that magnetic frustration destabilizes the low-temperature saturation of the AM phase. We attribute this behavior to thermally activated vortex motion, which introduces dissipation even before the onset of the Berezinskii–Kosterlitz–Thouless transition [21]. The observation of the finite slope confirms that the system remains thermally sensitive down to the base temperature, supporting an intrinsic AM phase at $f=0$ rather than an artifact of residual heating.
In summary, we have investigated voltage-controlled phase transitions in Josephson junction arrays patterned on an InAs/Al heterostructure. Using a dual-gate geometry, we find a robust anomalous metallic phase interrupting the superconductor-insulator transition, which can be smoothly tuned into a conventional metallic state with fully decoupled superconducting islands. At zero magnetic frustration, the anomalous metallic phase, characterized by vanishing low-temperature dependence, persists across more than three orders of magnitude in resistivity. A finite frustration suppresses the resistivity saturation, indicating thermally activated vortex motion and confirming thermal sensitivity down to the base temperature.

### v2

@@ Abstract
[…]
@@
Figure 3: (a) Resistivity, $\rho$, as a function of temperature, $T$, measured for different frame-gate voltage, $V_{\mathrm{FG}}$, values ranging from 1.060-1.060 to 1.225-1.225 V, at fixed $V_{\mathrm{TG}}=0$. The dashed black lines are linear fits of $\ln(\rho)$ versus $\ln(T)$ up to 7070 mK. (b) Low-$T$ resistivity slope, $d\ln(\rho)/d\ln(T)$, taken from the fits in (a), plotted against the average resistivity, $\bar{\rho}$, in the same $T$ range. Error bars are the standard uncertainties of the fitting.
@@
We note that $\rho$ at $V_{\mathrm{FG}}=-1.20$ V remains roughly $h/4e^{2}$ throughout the measured $T$ range [Fig. 3(a)]. For less negative $V_{\mathrm{FG}}$, $\rho$ drops below its normal-state value—in some cases by several orders of magnitude—before saturating at low $T$. For slightly more negative $V_{\mathrm{FG}}$, the first couple of traces still saturate, but at values higher than their high-$T$ resistivity. This behavior resembles a smooth crossover from a failed superconductor to a failed insulator within the anomalous metallic phase and may be attributed to quantum fluctuations [41, 34, 42]. We note that a similar phenomenology has been observed in cuprate superconductors [43], although the connection between underlying microscopic mechanisms remains unclear.
@@
Figure 4: (a) Average resistivity, $\bar{\rho}$, measured as a function of top-gate, $V_{\mathrm{TG}}$, and frame-gate, $V_{\mathrm{FG}}$, voltages, showing an enlargement of the region where the superconducting, insulating, and metallic states meet. The plot was constructed by averaging a set of $\rho$ maps over the temperature range from $T=15$ to 55 mK, taken every $2.5$ mK. (b) Low-$T$ $\rho$ slope, $d\ln(\rho)/d\ln(T)$, obtained by linear fitting of $\ln(\rho)$ versus $\ln(T)$ for the same data as in (a)
@@
All data discussed so far were taken at magnetic field carefully tuned to zero (for more details, see Supplemental Material [39]). We now turn to investigate the effect of an out-of-plane magnetic field, $B$, which introduces a superconducting phase gradient between neighboring islands [44]. To this end, we map the resistivity as a function of $T$ and magnetic frustration, $f=\Phi/\Phi_{0}$, where $\Phi=BA$ is the magnetic flux through a unit cell of area $A$ and $\Phi_{0}=h/2e$ is the superconducting flux quantum. We find that the critical temperature, $T_{\mathrm{C}}$, is maximal at $f=0$ and decays quickly as the array is frustrated; see Fig. 5(a). Additional peaks in $T_{\mathrm{C}}$ appear at several other frustration values, most prominently at $f=1/2$, corresponding to commensurate vortex configurations. While the periodicity of the frustration pattern is geometry independent, we expect that narrower junctions or increased island spacing—both reducing the inter-island coupling—would lead to a stronger suppression of superconductivity with frustration.
@@
In both cases, the emergence of a small but finite positive slope at nonzero $f$ indicates that magnetic frustration destabilizes the low-temperature saturation of the AM phase. We attribute this behavior to thermally activated vortex motion transverse to the current flow, generating a longitudinal electric field [45], which can introduce dissipation even before the onset of the Berezinskii–Kosterlitz–Thouless transition [21]. Additionally, the finite slope shows the system stays thermally sensitive at base temperature.
The fact that the resistance remains sensitive to $T$ down to the base temperature and evolves smoothly with gate voltage over a wide range suggests that the anomalous metallic phase observed at zero magnetic field is not a consequence of extrinsic effects, such as residual heating or environmental dissipation. Instead, the data are consistent with a scenario in which quantum fluctuations suppress long range phase coherence among proximitized superconducting islands. While the individual islands retain local Cooper pairing, the weakened inter-island coupling gives rise to superconducting phase fluctuations, leading to dissipation at low temperature. Possible sources of such fluctuations include microscopic inhomogeneities in Josephson coupling or charge offsets across the array, both of which may vary with gate voltage and enhance the effect of quantum fluctuations.
In summary, we have investigated voltage-controlled phase transitions in Josephson junction arrays patterned on an InAs/Al heterostructure. Using a dual-gate geometry, we find a robust anomalous metallic phase interrupting the superconductor-insulator transition, which can be smoothly tuned into a conventional metallic state with fully decoupled superconducting islands. At zero magnetic frustration, the anomalous metallic phase, characterized by vanishing low-temperature dependence, persists across more than three orders of magnitude in resistivity. A finite frustration suppresses the resistivity saturation, indicating thermally activated vortex motion and confirming thermal sensitivity down to the base temperature. These results provide new experimental insights into the origin of the anomalous metallic phase.
@@
Data availability— The data used to generate the figures in this work are available in Ref. [46].